Cougar park is shaped like a parallelogram and has an area of 1/16 square mile. It's base is 3/8 mile what is it's height?

Answer by Cromlix(4381)

Hi there,
Area of a parallelogram = height x base
1/16 = 3/8 x height
height = 1/16 divided by 3/8
height = 1/16 x 8/3
height = 1/6 mile
1/6 x 3/8 = 3/48 = 1/16 mile
Hope this helps :-)